Week of 050725

Open Actions from last week:

  • tape tests
  • testing of smoke tests
  • new FTS release / deployment plans.

On Shift: Maarten/Patricia

Monday:

Log:
  • Transfers fairly stable. Midnight Sat/Sun lost SARA (tech. problems with tape silo) Problems again last night with BNL. RAL 2-3 hours downtime.
  • Need to work out schedule with sites to debug / tune transfers.
  • FTS: 1% of transfers abort in gridftp. Grouped by site. Maarten will take a look...
  • FTS new version - still testing. Issue with MyProxy. Workaround... Need to agree / resolve release schedule.
  • James has populated ia32 disk servers. Olof to move to wan pool. (18 ia64 + 8-9 ia32)

New Actions:

Tuesday:

Log:
  • One problem with FTS006 - ASIS_OR_SPMA error - understood
  • LFC002 - GSSDATLAS Server - error in log due to list_replicax - fixed in 1.3.6
  • One of the new IA32 nodes still in closed state - Olof to check.
  • QF version of config service updated with config lockfile fix.

  • Wed/Thur - Olof/Sebastien/Ben in RAL. Jan on SMOD - will only turn up at 9.15

New Actions:

Wednesday

Log:
  • GRIDVIEW not stable - was an RGMA problem, but not sure if this explains all
  • LFC for LHCb deployed and LHCb gave OK for features
  • QF deployed already on FTS nodes (Gavin)
  • At PEB, experiments agreed to provide storage rates and capacity and num CPU for T0/1/2.

Actions:

  • Maarten: Look into GRIDVIEW numbers -
  • Jan/J-P: deploy new LFC on all pilot nodes DONE

Thursday

Log:
  • Still problems with Gridview - imported some of the outstanding data, RGMA team will import rest once system is stable
  • Waiting for IA64 version of RGMA - not urgent, have workaround (fri)
  • testing new version with load-gen to castorsrm

Actions:

  • Wait on Olof/... to return from RAL to reconfigure phedex nodes for DPM
  • discussion needed on resources for service phase at T1/T2. - for tuesday

Friday

Log:
  • Problem with lfc002 (atlas) - NO_CONTACT - seemed to not be a problem
  • tracking down problems in lcg-mon-gridftp and - r-gma - got a new version of the monitor that works with edg-rgma rpms
  • Andreas Unterkircher will work 50% on SC work until he joins in April
  • CMS have submitted test jobs they can do
  • Korean site for CMS joining the challenge. Need to start the process now.

Actions:

  • update LSF config and new version of lcg-mon-gridftp (Jan)
  • 'canned emails' for interventions (Patricia/Maarten)
  • start setting up Korean networking (David/James)

Todo:

  • new version of SRM (Ben/Vlado on tuesday)
  • QF for myproxy/oracle issue in FTS (install new RPM + restart)

Edit | Attach | Watch | Print version | History: r7 < r6 < r5 < r4 < r3 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r7 - 2007-02-02 - FlaviaDonno
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback